OpenStudy (anonymous):

help me correct my answers. tell me if i am wrong or right! WILL GIVE BEST RESPONSE if you help correctly :)

jigglypuff314 (jigglypuff314):

for the first, it is distributive property so 5x(x^4 - 3x^3 + 2x - 4) = (5x)(x^4) + (5x)(-3x^3) + (5x)(2x) + (5x)(-4) 5x^5 - 15x^4 + 10x^2 - 20x

jigglypuff314 (jigglypuff314):

the second is correct :)

jigglypuff314 (jigglypuff314):

if you understand foil (which is just an extended version of distributive property) then for the third (x^2 - 2) (x +3) = (x^2)(x) + (x^2)(3) + (-2)(x) + (-2)(3) x^3 + 3x^2 - 2x - 6

jigglypuff314 (jigglypuff314):

for the fourth... like how x^2 = x*x (x - 4)^2 = (x - 4)(x - 4) which you would foil

jigglypuff314 (jigglypuff314):

not really... try foiling it...

jigglypuff314 (jigglypuff314):

(x - 4)(x - 4) = (x)(x) + (x)(-4) + (-4)(x) + (-4)(-4) x^2 - 4x - 4x + 16 x^2 - 8x + 16 so it would be D

jigglypuff314 (jigglypuff314):

your welcome :)

jigglypuff314 (jigglypuff314):

(x+4)(-3x^2 - 4x + 2) distributive... (x)(-3x^2)+(x)(-4x)+(x)(2) + (4)(-3x^2)+(4)(-4x)+(4)(2)